OpenStudy (baby456):

14. A rectangle shown below has a perimeter of 40 inches. Its width is 5 inches? what is the area of a rectangle? PLease help medal and fan!

OpenStudy (baby456):



OpenStudy (anonymous):

Area = length x height. 5 x n = 40

OpenStudy (anonymous):

What number can you multiply times 5 to give you 40?

OpenStudy (baby456):

8

OpenStudy (anonymous):

that's your answer :)

OpenStudy (wolf1728):

Perimeter = 5 + 5 + length + length 40 = 10 + 2*length length = (40-10)/2 length = 15

OpenStudy (baby456):

ok

OpenStudy (baby456):

so the area is 5*15

OpenStudy (wolf1728):

I'd say so

OpenStudy (baby456):

so tje are is 75

OpenStudy (baby456):

could you pl;ease solve another probelm?

OpenStudy (wolf1728):

I guess so - time limit still in effect?

OpenStudy (baby456):

A rectangle shown below has a perimeter of 30 ft. its width is w feet.

OpenStudy (baby456):

what is the area of the rectangle?( the area should be expression in therms of variable w)

OpenStudy (baby456):



OpenStudy (wolf1728):

CORRECTION 2W + 2L = 30 (Perimeter) W + L = 15 Area = W * L Area = (15 - L) * L

OpenStudy (wolf1728):

Okay

OpenStudy (baby456):

Thank You!

OpenStudy (baby456):

I an going to close the question Now!

OpenStudy (wolf1728):

That's fine

OpenStudy (baby456):

thanks! bye!

OpenStudy (wolf1728):

c ya